# Symmetry Discovery Beyond Affine Transformations

### v1

@@ 4 Experiments
Our experiments in large part use simulated datasets. This appears to be common in current state of the art symmetry detection literature, as simulated data allows for direct comparison to ground truth symmetry. Our first experiment will compare our method to LieGAN in an effort to discover an affine symmetry. Our next experiment will demonstrate our method’s ability to identify symmetries which are not affine, thus advancing the current state of the art. Our third experiment will examine the estimation of an invariant feature of symmetry defined by a support vector classifier. Our final experiment before the appendices endeavors to show that symmetry-based feature engineering can improve model performance with regression. Additional auxiliary experiments are given in Appendix  C. Appendix  C.1 contains an experiment to estimate an invariant feature by first applying kernel density estimation on the (unlabeled) data. Appendix  C.2 applies level set estimation to a non-uniformly distributed circle embedded in $\mathbb{R}^{3}$. Appendix  C.3 detects discrete rotational symmetry on a dataset known to exhibit discrete symmetry [5]. Appendix  C.4 experiments with estimating the Riemannian metric tensor and discovering infinitesimal isometries. Finally, Appendix  C.5 applies level set estimation to an embedded submanifold of $\mathbb{R}^{10}$. All experiments were conducted on a personal computer using CPU. For repeated measures, the estimate is the mean, with error limits defined by one standard deviation.
@@ 4.1 Affine Symmetry Detection Comparison
Our first experiment compares the performance of our method to LieGAN, which is the current state of the art in symmetry detection. We will use simulated datasets where the “true” symmetries are known. Our first family of datasets consists of $N$ random samples generated from a 2-dimensional normal distribution with mean $(1,1)$ and covariance matrix $\left[\begin{smallmatrix}4&0\\ 0&1\end{smallmatrix}\right]$, where $N\in\{200,2000,20,000\}$. In each case, target function values are generated as $f=(x-1)^{2}+4(y-1)^{2}$ for each point $(x,y)$ in the dataset. The function exhibits symmetry: $X(f)=0$, with $X=-(4y-4)\partial_{x}+(x-1)\partial_{y}$ (or a scaling thereof). In the methodology employed by LieGAN, the ground truth infinitesimal generator is expressible in terms of a $3\times 3$ matrix, with generated transformations acting on the $z=1$ plane in three dimensions. Thus, in the case of LieGAN, the data is generated with mean $(1,1,1)$ and covariance given with the upper $2\times 2$ matrix as before, with all others values besides the bottom right value being 00, which bottom right value is $0.0001$ so as to approximately lie on the $z=1$ plane. Our own method uses the 2-dimensional data.
We will compare the results of our methods to the results of LieGAN using two metrics, namely cosine similarity to the ground truth and computation time. Our methods optimize in a 6-dimensional space, since a generic 2-dimensional vector field with affine component functions can be written as
$(a_{1}+a_{2}x+a_{3}y)\partial_{x}+(a_{4}+a_{5}x+a_{6}y)\partial_{y}.$
Thus, we will express the answer obtained by LieGAN as well as the ground truth in terms of our 6-dimensional space and compute the cosine similarity in the 6-dimensional space. The results are shown in Table 1 using 10 trials. These results suggest that if a sufficient amount of data is given, LieGAN is able to correctly identify the ground truth symmetry. Our method exhibits an advantage when less data is available, and our method takes less time, particularly when larger amounts of data are used. Our method also competes with LieGAN in terms of similarity to ground truth in higher sample sizes.
@@ 4.2 Non-Affine Symmetry Detection Comparison
After tuning the learning rate for learning $f=x^{3}-y^{2}$ as well as the learning rate for learning the estimated vector field, our method obtains a cosine similarity score of $0.9919\pm 0.0105$, having aggregated across 10 independent trials. Meanwhile, LieGAN obtains a cosine similarity score of $0.1945\pm 0.1133$ on 10 independent trials, being unable to discover symmetries of this complexity.
4.3 Invariant Features of a Model-Induced Class Probability Density
4.4 Symmetry-Assisted Regression
The symmetries of a distribution or shape on which data lies may correspond with another machine learning function such as target values in a regression problem. In this experiment, we will show that symmetry-based feature engineering can improve model performance on a regression task. The data for this experiment has been analyzed previously for the purpose of discovering discrete rotational symmetry [5], though we set $z=1$ for the purpose of demonstrating symmetry-informed feature engineering. We also search for discrete rotational symmetry for this dataset in Appendix  C.3.
${}f=\dfrac{1}{1+\arctan(y/x\text{ mod }2\pi/7)}.$ (15)
We estimate the invariant feature for this infinitesimal generator using our method. The flow parameter is approximately the polar angle, since the vector field given in Eq. 16 has a cosine similarity of $0.9994$ to the vector field corresponding to rotation in the plane about the origin. We thus transform the original dataset to coordinates $(\xi,\theta)$, where $\xi$ is the estimated invariant feature and where $\theta$ is the polar angle. We train a random forest model using the transformed data, and our results are given in Table 2. Based on the obtained $R^{2}$ scores in table 2, it appears that the regression task is more easily accomplished in the symmetry-based feature space than in the original feature space.
Dataset $R^{2}$ score (Train) $R^{2}$ score (Test)
Original 0.9488 0.5965
Transformed 0.9975 0.9965
@@ 5 Conclusion
We have also discussed potential novel applications of symmetry discovery, including model validation and symmetry-based feature engineering. Further potential applications are discussed in Appendices  A and  C.4. Our novel approach for detecting symmetry in data, the computational advantages of our method, and the demonstrated and potential applications of our method make our method of interest to the machine learning community.
@@ Appendix B Similar Methodology for Discrete Symmetries
For continuous symmetries, invariant quantities beyond those found to be of significance to the data set itself–the functions $h^{i}$–were used to define an “invariant feature space.” That is, a set of features were given such that when the data points were expressed in these features, any machine learning task performed on this new feature space would necessarily be invariant with respect to the given symmetry group. For discrete symmetry detection, the transformations can be given explicitly, so that existing methods for building invariance into a model that require the symmetry group to be given beforehand can be employed. However, it may also be possible–perhaps depending on the discovered symmetries–to calculate features $h^{i}$ that are invariant with respect to the discovered discrete symmetries. For example, the function $h^{1}=y$ is invariant to reflection about the $x=0$ line. The identification of features which are invariant with respect to a fixed collection of symmetries would allow one to transform points from the original feature space to the invariant feature space, so that any machine learning task would be invariant with respect to the given symmetry group.
Appendix C Additional experiments
C.1 Invariant Feature Construction
@@
with $\sum_{i=1}^{6}a_{i}^{2}=1$. After estimating $h$, we plot a grid of function values of $h$. The result is pictured alongside the visualization of the estimated probability distribution of the data (Figs. 4 and 4).
@@ C.2 Level Set Estimation
We wish to continue looking for suitable functions that, when combined with 18, describe the manifold on which the data lies. Thus, we will expand our search space to quadratic functions. Before doing so, in an attempt to avoid discovering equations degenerate with respect to 18, we will project the 3-dimensional data onto the space defined by our discovered affine functions–in this case, the approximate $(x,y)$ plane (though technically $z=1$). A plot of the result is obtained in 6. We then apply 10 on the reduced space, now with quadratic features. The discovered component is given as
@@
In the search for infinitesimal generators which annihilate both 18 and 19 in accordance with 13, we perform four (relatively fast) searches: first, among vector fields with constant components; second among vector fields with linear components (no constant terms); third, among vector fields with affine components; fourth among vector fields with quadratic components. In this case, since the effective dimension of the dataset is 1, we limit our searches to those where the number of columns of $W$ is 1. Our best result is obtained in the second case, and the estimated vector field is
@@
to which the reader can compare with $x^{2}+y^{2}$. A plot of the function values of $f$ is given in 6.
@@
Figure 6: A visual representation of the function values of the estimated invariant function of $X$ as given in 20.
It appears that Figure 6 is more faithful to the ground-truth invariant function than 4. We do not believe this indicates the inferiority of symmetry detection using probability density estimation. Rather, we believe this may have been caused by the estimation of gradients of the probability distribution: the gradients of the level set functions were computed exactly. This may indicate that a parametric approach to density estimation may allow for better approximation of gradients.
C.3 Discrete Rotation
@@
where $P$ is the vectorization of $p$. We choose $\theta>\pi/6$ so as to find transformations away from the trivial $\theta=0$. Our optimization, performed using python’s scipy package, yields a result of $\theta\approx 0.8972$, which is $\frac{2\pi}{7}$ to four decimal places, and our cosine similarity to ground truth is $1.0000$. The total computation time, including the estimation of the density, is approximately 48 seconds.
We repeat the experiment using only 1000 data points, recovering $\theta\approx 0.8333$, in approximately $0.2$ seconds. This yields a cosine similarity to ground truth of $0.9979$. On the other hand, when the amount of data for the experiment with LieGAN is moderately reduced, the model performance suffers greatly. The intended generator is
@@
The cosine similarity of these two generators is $0.1956$. Thus, our approach seems to have greater potential to handle discrete symmetry with smaller datasets in two dimensions, in addition to having a friendlier interface.
C.4 Infinitesimal Isometries
@@
The experiment begins under the assumption that the starting data set is $\mathcal{D}$, and that the data set $\tilde{\mathcal{D}}$ has been produced from $\mathcal{D}$ by means of a dimensionality reduction algorithm. The computation begins by learning functions $x(u,v,w)$, $y(u,v,w)$, $z(u,v,w)$, and $t(u,v,w)$ with polynomial regression. This we do, with results varying only, it seems, due to numerical approximation error, so that we take 27 to be our estimation.
We will need to compute the Killing vectors for the 3-dimensional space, for which we will need the Riemannian metric. We assume that the metric in the original 4-dimensional space is the Euclidean metric, and we use 27 to pull back this Euclidean metric. The result can be expressed as the following matrix in the coordinates $(u,v,w)$:
@@
This is equivalent to the formulation given in 13 with the feature matrix being defined by the components of the known Killing vectors and with $W$ having six rows and a single column. We handle this particular task using McTorch [15], which is a manifold optimization library compatible with pytorch. The result is a cosine similarity score of $1.0000$ with respect to the ground truth of
@@
C.5 10-Dimensional Reduction using Level Set Estimation
@@
The equations given in 34 could be used to define a function $\Theta:\mathbb{R}^{4}\to\mathbb{R}^{10}$, allowing a Riemannian metric tensor to be given on $\mathbb{R}^{4}$ by pulling back the Euclidean metric tensor on $\mathbb{R}^{10}$, much like what we demonstrated in a previous example. However, our goal here is level set estimation. Our dataset, by construction, admits two different and equally valid level set descriptions. The first is given by the following equations:

### v2

@@ 4 Experiments
Our first experiment will compare our method to LieGAN in an effort to discover an affine symmetry. Our next experiment will demonstrate our method’s ability to identify symmetries which are not affine. The third experiment demonstrates that symmetry can occur in real data, and that our method can detect approximate symmetry in real data. Our last main experiment deals with pre-determined models which are not polynomials–in particular, when the ground truth symmetry is not expressible in terms of polynomial coefficient functions. Additional experiments appear in Appendix D: some highlights are symmetry of a model-induced probability distribution, discrete symmetry estimation, and estimation of infinitesimal isometries.
@@ 4.1 Affine Symmetry Detection Comparison
Our first experiment compares the performance of our method to LieGAN, which is the current state of the art in symmetry detection. We use simulated data where the “true” symmetries are known. Our three datasets for this experiment consist of $N$ random samples generated from a 2-dimensional normal distribution with mean $(1,1)$ and covariance matrix $\left[\begin{smallmatrix}4&0\\ 0&1\end{smallmatrix}\right]$, where $N\in\{200,2000,20,000\}$. In each case, target function values are generated as $f=(x-1)^{2}+4(y-1)^{2}$ for each point $(x,y)$ in the dataset. The function exhibits symmetry: $X(f)=0$, with $X=-(4y-4)\partial_{x}+(x-1)\partial_{y}$ (or a scaling thereof). In the methodology employed by LieGAN, the ground truth infinitesimal generator is expressible in terms of a $3\times 3$ matrix, with generated transformations acting on the $z=1$ plane in three dimensions. Thus, in the case of LieGAN, the data is generated with mean $(1,1,1)$ and covariance given with the upper $2\times 2$ matrix as before, with all other values besides the bottom right value being 00, which bottom right value is $0.0001$ so as to approximately lie on the $z=1$ plane. Our own method uses the 2-dimensional data.
[…]
We compare the results of our method to the results of LieGAN using two metrics, namely similarity to the ground truth defined by Equation (15) and computation time. The results are shown in Table 1 and use 10 trials for each $N$. These results suggest that if a sufficient amount of data is given, LieGAN is able to correctly identify the ground truth symmetry. However, our method exhibits a clear advantage for $N=200$ while offering comparable similarity scores to LieGAN for larger sample sizes. Perhaps the most striking feature of Table 1, however, is the dramatic difference in time needed to correctly identify the symmetry. Our method offers a clear advantage in terms of computation time as the number of samples increases.
@@ 4.2 Non-Affine Symmetry Detection Comparison
We first use polynomial regression to approximate $f=x^{3}-y^{2}$. Then, we estimate $W$ in Equation (13) using the $L_{1}$ loss function, selecting the Riemannian Adagrad optimization algorithm [16] with a learning rate of $0.1$, training for $5000$ epochs. Our method then obtains a similarity score of $0.9983\pm 0.0006$, having aggregated across 10 independent trials. Meanwhile, LieGAN obtains a similarity score of $0.0340\pm 0.0142$ on 10 independent trials, being unable to discover symmetries of this complexity.
4.3 Symmetry Detection using Realistic Data
Our next experiment uses real data that is publicly available, which dataset we refer to as the “Bear Lake Weather” dataset [17]. The dataset gives daily weather attributes. It contains $14,610$ entries with 81 numeric attributes including the daily elevation level of the lake. The dataset contains precisely 40 years’ worth of data from October of 1981 through September of 2021.
We believe an understanding of the behavior of the weather in the past is relevant to this problem. Therefore, we first construct time series of length in dimensions by means of a sliding window of length days: the first time series is constructed using the first 1461 days (the number of days in four years). The next time series is constructed using the second day through day 1462, and so forth.
After converting the raw data to time series data, we apply a transformation on the data meant to extract time-relevant features of the data known as the Multirocket transform [18]. We select 168 kernels in our transform. The Multirocket transform transforms the data from $13,149$ time series of length $1461$ in $81$ variables to tabular data: $13,149$ entries in 1344 variables.
For such high-dimensional data, we turn to PHATE, a state of the art data visualization method [19]. Using PHATE, we reduce the dimension to $2$, so that our new dataset has $13,149$ entries in 2 variables. The resulting data appears to approximately lie on a circular shape and is shown in Figure 1.
Figure 1 suggests that the data is largely periodic. In fact, further experimentation reveals that the points for a given calendar year correspond to a traversal around the circular-like shape. Thus, for the analysis of non-seasonal weather patterns, it may be of use to examine features of this embedded dataset which are invariant under the periodic transformation, the approximate symmetry.
Figure 1: The ROCKET-PHATE embedded Bear Lake Weather dataset, colored by invariant function value.
Indeed, our method reveals an approximate invariant function given by
$00.33592x^{2}+0.94189y^{2}-2.9743\cdot 10^{-4}=$ 0 .
This result was obtained using level set estimation, assuming a function expressible as $ax^{2}+by^{2}+c$, using the Mean Square Error loss function with Equation (10), optimized using Riemannian stochastic gradient descent with a learning rate of $0.001$ and trained for $5000$ epochs. This experiment shows that symmetry can occur in real data, and that our method can detect symmetry and estimate invariant functions for real data.
4.4 Symmetry Detection with Non-Polynomial Ground Truth
This experiment deals with a case in which the ground truths are not strictly polynomial functions. We generate $2048$ numbers $x_{i}$ and $2048$ numbers $y_{j}$ each from $U(0,2\pi)$. Next, for each pair $(x_{i},y_{i})$, we obtain $z_{i}$ by means of $z_{i}=\sin(x_{i})-\cos(y_{i})$, so that a ground-truth level set description of the data is given as $z-\sin(x)+\cos(y)=0.$
We first apply our level set estimation method to estimate this level set. We optimize the coefficients of the model
$a_{0}+a_{1}x+a_{2}y+a_{3}z+a_{4}\cos(x)+a_{5}\cos(y)+a_{6}\cos(z)+a_{7}\sin(x)+a_{8}\sin(y)+a_{9}\sin(z)=0$
subject to $\sum_{i=0}^{9}a_{i}^{2}=1$. In light of Equation 10, the matrix $B$ has a row for each of the 2048 tuples $(x_{i},y_{i},z_{i})$, and 10 columns, which columns correspond to the 10 different feature functions in our pre-determined model. The vector $w$ contains all 10 parameters $\{a_{i}\}_{i=0}^{9}$.
Using the $L_{1}$ loss function and the Riemannian Adagrad optimization algorithm [16] with learning rate $0.01$, our estimated level set description is
$-0.57737z-0.57713\cos(y)+0.57756\sin(x)=0,$
which is approximately equivalent to the ground truth answer up to a scaling factor.
Having demonstrated that our method easily extends to non-polynomial pre-determined features, we now wish to examine a case in which a polynomial pre-determined model of symmetry does not contain the ground truth vector field. For this, we use the same dataset, albeit we define $f_{i}=z_{i}$, so that we are approximating symmetries of $f(x,y)=\sin(x)-\cos(y)$. A ground truth vector field which characterizes the symmetry of $f$ is given as
$X=\sin(y)\partial_{x}-\cos(x)\partial_{y}.$
Applying our method with a pre-determined model of degree 2 polynomial coefficients gives an estimated (using Equation 13) vector field $\hat{X}$ of
$\hat{X}=\left(0.7024-0.1874x-0.2203y+0.0121x^{2}+0.0242xy+0.0133y^{2}\right)\partial_{x}$
$+\left(-0.5783+0.2665x+0.1236y-0.0311x^{2}-0.0150xy-0.0097y^{2}\right)\partial_{y}.$
This result was obtained using the $L_{1}$ loss function, the Riemannian Adagrad optimizer [16] with learning rate $0.1$, training for $5000$ epochs. We also note that sim$(X,\hat{X})=0.62$, using the similarity score defined in Equation (15). This demonstrates that our method can be applied even when a ground truth vector field is not covered by the vector field pre-determined model.
@@ 5 Conclusion
We have also discussed potential novel applications of symmetry discovery, including model validation and symmetry-based feature engineering. Further potential applications are discussed in Appendices  A and  D.4. Our novel approach for detecting symmetry in data, the computational advantages of our method, and the demonstrated and potential applications of our method make our method of interest to the machine learning community.
@@ Appendix B Similar Methodology for Discrete Symmetries
To estimate discrete symmetries, we assume that a symmetry belongs to a parametric family of transformations, which parameters we denote by $\mathbf{w}$. We define, for each $x_{i}$ in a given dataset,
${}\hat{y}_{i}=S(f(x_{i});\mathbf{w})-f(x_{i}),\qquad y_{i}=0,$ (16)
and we seek to minimize $L(\hat{y}_{i},y_{i})$ for a smooth loss function $L$, typically by means of constrained optimization. For example, we consider the 2-dimensional task of determining a line of reflection about which a function $f(x,y)$ is symmetric. The reflection of a point $(x,y)$ about a line in the form of $ax+by=0$ can be written as
${}S(x,y;a,b)=\dfrac{1}{a^{2}+b^{2}}\begin{bmatrix}a^{2}-b^{2}&-2ab\\ -2ab&b^{2}-a^{2}\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}.$ (17)
Minimizing $L(\hat{y}_{i},y_{i})$ for a smooth loss function then estimates the parameters $(a,b)$ which characterize the line for which $f(x,y)$ is approximately reflectively symmetric. Since the line $ax+by=0$ is equivalent to $cax+cby=0$ for some constant $c$, our method is best accomplished by means of constrained optimization, which, in this example, is easily accomplished by means of $a^{2}+b^{2}=1$.
@@
For continuous symmetries, invariant quantities beyond those found to be of significance to the data set itself–the functions $h^{i}$–were used to define an “invariant feature space.” That is, a set of features were given such that when the data points were expressed in terms of these features, any machine learning task performed on this new feature space would necessarily be invariant with respect to the given symmetry group. For discrete symmetry detection, the transformations can be given explicitly, so that existing methods for building invariance into a model that require the symmetry group to be given beforehand can be employed. However, it may also be possible–perhaps depending on the discovered symmetries–to calculate features $h^{i}$ that are invariant with respect to the discovered discrete symmetries. For example, the function $h^{1}=y$ is invariant to reflection about the $x=0$ line. The identification of features which are invariant with respect to a fixed collection of symmetries would allow one to transform points from the original feature space to the invariant feature space, so that any machine learning task would be invariant with respect to the given symmetry group.
We also note that our assumption that the discrete symmetries sought originate in parametric families of transformations is limiting. In particular, despite previous work conducted with permutation groups [6], it does not appear that our method can be applied to permutation groups.
Appendix C Limitations
Our method is an improvement to state of the art methods for continuous symmetry detection, both for computational reasons and due to the fact that our method can detect continuous symmetries which are not affine transformations. However, there are limitations to our approach. First, as we have mentioned previously, is the issue which arises when vector fields $X$ and $fX$ are both present in the search space. We have suggested that a symbolic approach may resolve this, though non-symbolic approaches such as ours must grapple with this issue. Our approach of searching for polynomials of lower order first partially resolves this issue. Another promising resolution is to restrict the search space of vector fields to particular types of symmetries, such as isometries, which is experimented with in Appendix D.4.
Another limitation is that our method requires an assumption about the coefficient functions of the vector fields. Herein, we primarily use polynomial functions, and although polynomials are universal approximators [23], the number of polynomial terms for a given coefficient function can increase rapidly, particularly when the number of dimensions is high. We believe this limitation is best resolved, as with the previous limitation, by restricting the search space to particular types of symmetries.
Though our method is comparatively efficient, the reliance on the elbow curve to determine both the number of discovered symmetries and, when level set estimation is used, the number of coefficients in the level set, is a limitation. This limitation is of a more persistent nature than the previous limitations, appearing in analogous form as the dimension hyperparameter in manifold learning methods, or as the number of clusters in $k$-means clustering. It is possible that our method would be improved by another means of selecting this number, however.
We also mention the limitations inherited via the choice of optimization of equations (13) and (14). Cost function selection, optimization algorithm selection, hyperparameter selection, together with the potential for poor parameter initialization all constitute limitations to our method. We believe the optimization task is much more straight-forward than training current state of the art symmetry detection models, however, as other methods require fine-tuning neural networks.
Lastly, our notion of similarity of vector fields, while adequate for our experiments, is limited. Primarily, when a collection of vector fields is estimated and taken to constitute the infinitesimal generators for the symmetry group, it is not clear how to compare the estimated collection of vector fields to a ground truth collection. The similarity score also depends on the range of a given dataset, and the aggregation of component-wise similarities via the arithmetic mean is naïve, leaving room for further refinement. In particular, in the presence of a metric tensor $g$, one may seek to define the similarity score by means of $\mathbb{E}\left[\langle X,\hat{X}\rangle_{g}\right]$, where the inner product $\langle X,\hat{X}\rangle_{g}$ is defined by the metric $g$ on the tangent space of the manifold, and where the expected value is computed by means of the underlying distribution of the data. However, exploration of such refinements are left as future work.
Appendix D Additional experiments
D.1 Invariant Feature Construction
@@
with $\sum_{i=1}^{6}a_{i}^{2}=1$. After estimating $h$, we plot a grid of function values of $h$. The result is pictured alongside the visualization of the estimated probability distribution of the data (Figs. 3 and 3).
@@ D.2 Level Set Estimation
We wish to continue looking for suitable functions that, when combined with 19, describe the manifold on which the data lies. Thus, we will expand our search space to quadratic functions. Before doing so, in an attempt to avoid discovering equations degenerate with respect to 19, we will project the 3-dimensional data onto the space defined by our discovered affine functions–in this case, the approximate $(x,y)$ plane (though technically $z=1$). A plot of the result is obtained in 5. We then apply 10 on the reduced space, now with quadratic features. The discovered component is given as
@@
In the search for infinitesimal generators which annihilate both 19 and 20 in accordance with 13, we perform four (relatively fast) searches: first, among vector fields with constant components; second among vector fields with linear components (no constant terms); third, among vector fields with affine components; fourth among vector fields with quadratic components. In this case, since the effective dimension of the dataset is 1, we limit our searches to those where the number of columns of $W$ is 1. Our best result is obtained in the second case, and the estimated vector field is
@@
to which the reader can compare with $x^{2}+y^{2}$. A plot of the function values of $f$ is given in 5.
@@
Figure 5: A visual representation of the function values of the estimated invariant function of $X$ as given in 21.
It appears that Figure 5 is more faithful to the ground-truth invariant function than 3. We do not believe this indicates the inferiority of symmetry detection using probability density estimation. Rather, we believe this may have been caused by the estimation of gradients of the probability distribution: the gradients of the level set functions were computed exactly. This may indicate that a parametric approach to density estimation may allow for better approximation of gradients.
D.3 Discrete Rotation
@@
where $P$ is the vectorization of $p$. We choose $\theta>\pi/6$ so as to find transformations away from the trivial $\theta=0$. Our optimization, performed using python’s scipy package, yields a result of $\theta\approx 0.8972$, which is $\frac{2\pi}{7}$ to four decimal places, and our similarity to ground truth is $1.0000$. The total computation time, including the estimation of the density, is approximately 48 seconds.
We repeat the experiment using only 1000 data points, recovering $\theta\approx 0.8333$, in approximately $0.2$ seconds. This yields a similarity to ground truth of $0.9979$. On the other hand, when the amount of data for the experiment with LieGAN is moderately reduced, the model performance suffers greatly. The intended generator is
@@
The similarity of these two generators is $0.1956$. Thus, our approach seems to have greater potential to handle discrete symmetry with smaller datasets in two dimensions, in addition to having a friendlier interface.
D.4 Infinitesimal Isometries
@@
The experiment begins under the assumption that the starting data set is $\mathcal{D}$, and that the data set $\tilde{\mathcal{D}}$ has been produced from $\mathcal{D}$ by means of a dimensionality reduction algorithm. The computation begins by learning functions $x(u,v,w)$, $y(u,v,w)$, $z(u,v,w)$, and $t(u,v,w)$ with polynomial regression. This we do, with results varying only, it seems, due to numerical approximation error, so that we take 28 to be our estimation.
We will need to compute the Killing vectors for the 3-dimensional space, for which we will need the Riemannian metric. We assume that the metric in the original 4-dimensional space is the Euclidean metric, and we use 28 to pull back this Euclidean metric. The result can be expressed as the following matrix in the coordinates $(u,v,w)$:
@@
This is equivalent to the formulation given in 13 with the feature matrix being defined by the components of the known Killing vectors and with $W$ having six rows and a single column. We handle this particular task using McTorch [16], which is a manifold optimization library compatible with pytorch. The result is a similarity score of $1.0000$ with respect to the ground truth of
@@
D.5 10-Dimensional Reduction using Level Set Estimation
@@
The equations given in 35 could be used to define a function $\Theta:\mathbb{R}^{4}\to\mathbb{R}^{10}$, allowing a Riemannian metric tensor to be given on $\mathbb{R}^{4}$ by pulling back the Euclidean metric tensor on $\mathbb{R}^{10}$, much like what we demonstrated in a previous example. However, our goal here is level set estimation. Our dataset, by construction, admits two different and equally valid level set descriptions. The first is given by the following equations:
@@
D.6 Invariant Features of a Model-Induced Class Probability Density
D.7 Symmetry-Assisted Regression
The symmetries of a distribution or shape on which data lies may correspond with another machine learning function such as target values in a regression problem. In this experiment, we will show that symmetry-based feature engineering can improve model performance on a regression task. The data for this experiment has been analyzed previously for the purpose of discovering discrete rotational symmetry [5], though we set $z=1$ for the purpose of demonstrating symmetry-informed feature engineering. We also search for discrete rotational symmetry for this dataset in Appendix  D.3.
${}f=\dfrac{1}{1+\arctan(y/x\text{ mod }2\pi/7)}.$ (38)
We estimate the invariant feature for this infinitesimal generator using our method. The flow parameter is approximately the polar angle, since the vector field given in Eq. 39 has a similarity of $0.9994$ to the vector field corresponding to rotation in the plane about the origin. We thus transform the original dataset to coordinates $(\xi,\theta)$, where $\xi$ is the estimated invariant feature and where $\theta$ is the polar angle. We train a random forest model using the transformed data, and our results are given in Table 2. Based on the obtained $R^{2}$ scores in table 2, it appears that the regression task is more easily accomplished in the symmetry-based feature space than in the original feature space.
Dataset $R^{2}$ score (Train) $R^{2}$ score (Test)
Original 0.9488 0.5965
Transformed 0.9975 0.9965
Appendix E A Note about the Freedom Present in Identifying Vector Fields
Given a vector field $X$ and a vector field $fX$, a function $h$ is $X$-invariant if and only if it is $fX$-invariant. If $h$ is $X$-invariant, the flow of $X$ is a symmetry of $h$, so that the flow of $fX$ is also a symmetry of $h$: this follows from our discussion of vector fields in section 2.
For example, a circle centered at the origin, characterized by $F=0$ where $F(x,y)=x^{2}+y^{2}-r^{2}$ for some real number $r$, exhibits rotational symmetry described by $X=-y\partial_{x}+x\partial_{y}$. However, $\frac{1}{y}X(F)=0$, so that the flow of the vector field $-\partial_{x}+\frac{x}{y}\partial_{y}$, where defined, is also a symmetry of the circle.
This may seem to introduce a theoretical problem requiring greater care when detecting non-affine symmetry. However, our proposed method of constructing models which are invariant with respect to the symmetry group requires only the identification of functions which are invariant with respect to the vector fields. Thus, both $X$ and $fX$ are valid answers for “ground truth” symmetries, since a function $h$ is $X$-invariant if and only if it is $fX$-invariant.
We note that this lack of uniqueness does present a challenge when estimating suitable vector fields. When estimating the symmetries via constrained regression, $X$ and $fX$ may or may not both be present in the search space. As mentioned in section 3, one potential workaround is owing to the fact that symmetry estimation can be done symbolically. Consider the example of $F=0$ given previously in this section. The Jacobian matrix of $F$ is given as
$J=\begin{bmatrix}2x&2y&2z\end{bmatrix},$
and a basis for the nullspace of $J$ (using functions as scalars) is
$\left\{\begin{bmatrix}-y\\ x\\ 0\end{bmatrix},\begin{bmatrix}0\\ -z\\ y\end{bmatrix}\right\},$
which vectors correspond to tangent vectors $-y\partial_{x}+x\partial_{y}$ and $-z\partial_{y}+y\partial_{z}$, respectively. The flows of these vector fields correspond with rotations (about the origin) in the $(x,y)$ and $(y,z)$ planes, respectively. However, only two vector fields can be recovered using this method, which in this case has neglected another symmetry, namely $-z\partial_{x}+x\partial_{z}$, corresponding to rotations in the $(x,z)$ plane. In fact, each of the three vector fields corresponding to rotations about a coordinate axis are isometries of the sphere described by $F=0$. We should note, however, that the generator for rotations in the $(x,z)$ plane can be expressed as a linear combination of the other rotations using functions as scalars:
$\dfrac{z}{y}\left(-y\partial_{x}+x\partial_{y}\right)+\dfrac{x}{y}\left(-z\partial_{y}+y\partial_{z}\right)=-z\partial_{x}+x\partial_{z}.$
Therefore, it is plausible that applications exist in which a symbolic approach to continuous symmetry discovery may be sufficient. Our approach does not rely on symbolic software, however.
We also mention that the issue of unwanted freedom can be addressed by identifying “special” vector fields and reducing the search space of vector fields to linear combinations of these special vector fields known as Killing vectors. This would eliminate the uniqueness problem and is experimented with in Appendix D.4. This approach is a special case of symmetry detection where only isometries can be detected.
Another point which may help to address concerns about our handling of the difficult nature of non-affine symmetry detection is in the relationship between the flows of $X$ and $fX$ generally. The trace of the flow of $X$ through the point $p$ is characterized by the level set $h_{i}=c_{i}$, where $\{h_{i}\}$ is a complete set of $X$-invariant functions. Since a function is $X$-invariant if and only if it is $fX$-invariant, a complete set of invariant functions for $fX$ can be taken to be $\{h_{i}\}$ without loss of generality. Thus, the level set $h_{i}=c_{i}$ also characterizes the trace of the flow of $fX$ through the point $p$. This argument assumes that the flows of $X$ and $fX$, as well as the vector fields themselves, are well-defined in an open neighborhood about the point $p$.
In fact, in our example above with $F=x^{2}+y^{2}-r^{2}$, the flow of $-\partial_{x}+\frac{x}{y}\partial_{y}$, assuming $y>0$, is given as
$\Phi(t,(x,y))=\left(x-t,\sqrt{y^{2}+2tx-t^{2}}\right),$
which trace is a (part of a) circle (we note that the flow parameter is $-x$). The trace of this flow through a point is equivalent to the trace of $X$ through the same point.